log x/n=log(x^1/n)
Answer by mukhopadhyay(490)

log x/n=log(x^1/n)
=> log x - log n = 1/n log x
=> log x - 1/n log x = log n
=> log x (1-1/n) = log n
=> log x = [n/(n-1)]*log n
=> x = 10^[(n/n-1)log n]